given the x-intercept for the grap of a quadratic function are 3 and -2, write a possible formula for the quadratic function, in factored form

Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

given the x-intercept for the grap of a quadratic function are 3 and -2, write a possible formula for the quadratic function, in factored form
----------
With x-ints of a & b (a,0) & (b,0), an equation is
(x-a)*(x-b) = 0

Answer by Theo(13342)

the zeroes of the quadratic are at x = 3 and x = -2
subtract 3 from the first equation to get x - 3 = 0
add 2 to the second equation to get x + 2 = 0
those are your factors of the quadratic.
you get (x-3) * (x+2) = 0
simplify to get x^2 + 2x - 3x - 6 = 0
combine like terms to get x^2 -x - 6 = 0
here's what the factors and the equation look like on a graph.

$$$

both equations draw the same graph because they are equivalent.